OpenStudy (anonymous):

How do you write 1.9 × 104 in standard form? FAN ME

OpenStudy (campbell_st):

I'll presume the question is \[1.9 \times 10^4\] you need to know \[10^4 = 10 000\] so you need to find the value of \[1.9 \times 10000 = ?\]

OpenStudy (anonymous):

whats up